OpenStudy (anonymous):

The graph of which of the following will be parallel to the graph of y = 2x – 4? 2x + y = 4 (i think this one is right) 4x - 2y = -12 y = -4x + 8 y = 3x + 4

OpenStudy (cwrw238):

the slope of graph = 2 the correct option will also have slope of 2

OpenStudy (anonymous):

So was I right? Since the original one has 2x and so did the answer I chose? because in ' y = mx + b' the 'm' stands for slope? They both have 2?

OpenStudy (cwrw238):

no that one is not right rearrangge it to the form y = mx + b you get y = -2x + 4 which has a slope of -2 not 2

OpenStudy (cwrw238):

you need to rearrange the functions to the form y = mx + b so you can compare them with the function in the question

OpenStudy (cwrw238):

the last 2 cannot be parallel as they have slopes of -4 and 3

OpenStudy (anonymous):

So would it be the second one? If you divide 4 by two it would be 2? I already eliminated the last option.

OpenStudy (cwrw238):

convert the second one to y = mx + b form and see what you get

OpenStudy (cwrw238):

4x - 2y = 12 4x - 12 = 2y 2y = 4x - 12 y = 2x - 6 yes